For the following probability distribution μ =
xi| 250 300 400
pi| 0.30 0.50 0.20

a. 0.50
b. 300
c. 305
d. 317
: For the following probability distribution μ =
xi| 250 300 400
pi| 0.30 0.50 0.20

a. 0.50
b. 300
c. 305
d. 317

Answer by Edwin McCravy(2087)

For the following probability distribution μ = 
xi|   250    300   400
pi|  0.30   0.50  0.20

mu=sum((x[i]p[i]),i=1,n)=x[1]p[1]+x[2]p[2]+x[3]p[3]=250*0.30+300*0.50+400*0.20)=75+150+80=305

Edwin